# On the Verification of Control Flow Attestation Evidence

### v1

@@ 4.4 Patch Generator
4.4.1 Patching Use-After-Free
@@
4.4.2 Patching Buffer Overflows
@@
    Replace the original call to the function from [T3] with a trampoline to the patched version.
Detecting Buffer Bounds
Buffer overflows are eliminated via bound checks before memory accesses according to the buffer size. From the previous phase, the Exploit Locator identified the offending instruction to be located at $addr_{Acc}$ and that it causes an overflow onto $base$. This suggests that a buffer exists next to $base$ that must have been overwritten by a loop before the overflow occurs. It also implies that the offending instruction overflowing to $base$ is the same as that used to write data to the buffer.
@@
As depicted in Fig. 7, the buffer’s definition is influenced by -4(r4) at 0xe27e, which is in turn initialized by r11 at 0xe206. Tracking of definitions continues in this manner until a known pointer value is reached, as described in Sec. 4.3.1. In this example, definitions are tracked until 0xe0e4 since the stack pointer (a known pointer value) has been reached. Thus, Exploit Locator saves 0xe0e4 as its initialization point, as $addr_{lower}$ (the address defining the lower bound of the exploited buffer) and saves the value of $reg_{Acc}$ (sp in this example) as the lower bound of the overflown buffer at this instruction.
@@
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects in addition to the buffer of interest), narrowing the scope of the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by limiting buffer overflows into the buffer’s own stack frame.
Recording the Bounds
After identifying $addr_{lower}$ and $addr_{upper}$, SABRE completes T1 and has estimated the bounds of the buffer that overflows into $addr_{target}$. To complete T2, the values of $reg_{Acc}$ must be saved at both locations. This is completed by placing a trampoline both at $addr_{lower}$ and $addr_{upper}$ to preserve their values in a dedicated register. Fig. 9 depicts the entire patch for buffer overflow vulnerabilities and shows how these instructions are added for recording the bounds.
Figure 9: Example patch for buffer overflow vulnerabilities.
@@
Inserting the Bounds Check
Then, Patch Generator proceeds to patch the offending memory instruction by prepending it with instructions that check whether the address being accessed is outside the valid boundary. If it is, the offending instruction is skipped. For the patch depicted in Fig. 9, the prepended instructions are highlighted in red and correspond to:
@@
Calling the Safe Function
To realize T4, Patch Generator rewrites the original call instruction that leads to the function containing the buffer overflow. It changes this call instruction to instead call the patched version. This is depicted in Fig. 9 by the instructions highlighted in green.
@@ 4.5 Patch Validator
Before deploying the patched software on $\sf{\mathcal{P}rv}$, it must be validated to ensure its effectiveness against the previously identified attack. To achieve this, this module first builds an updated CFG of the patched binary. Then, it remaps the destination addresses in $\mathcal{C}F_{Slice}$ to their respective addresses in the updated CFG. It also inserts new entries into $\mathcal{C}F_{Slice}$ as necessary to account for the patching behaviors (e.g., a trampoline to/from the patch). Patch Validator then performs symbolic data flow analysis over the translated $\mathcal{C}F_{Slice}$, similar to how it is done in Sec. 4.3. When the analysis reaches the end of $\mathcal{C}F_{Slice}$ without $addr_{target}$ being corrupted, Patch Validator considers this patch effective. When the corruption still occurs, the attack was likely caused by another vulnerability. In this case, SABRE generates a report indicating to $\sf{\mathcal{V}rf}$ that further manual analysis is required.
(a) Static Analysis: Build CFG
(b) Path Verifier: Verify $\mathcal{C}F_{Log}$
(c) Exploit Locator: Backwards Traversal
(d) Exploit Locator: Symbolic DF Analysis
(e) Patch Generator: Generate Patch
(f) Patch Generator: Update ELF
(g) Patch Validator: Translate $\mathcal{C}F_{Log}$
(h) Patch Validator: Build Patched CFG
(i) Patch Validator: Symbolic DF Analysis
Figure 10: SABRE submodule timing for detecting and patching crafted control flow attacks inserted into a set of BEEBs [67] MCU software. This figure compares detecting/patching times for buffer overflow (OVF) and use-after-free (UAF) vulnerabilities across two MCUs and CFA architectures: hardware-based CFA atop MSP430 [6] and TEE-based CFA atop ARM Cortex-M33 [10].
5 Implementation details
@@
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. For the evaluation, we use applications from the BEEBs benchmark for embedded platforms [67]. These applications were compiled for MSP430 and ARM Cortex-M and executed atop unmodified CFA architectures, namely ACFA [6] and TRACES [10], to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. We believe that SABRE concepts generalize to several CFA architectures and select ACFA and TRACES as case studies due to their open-source availability.
@@
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy [68] for evaluating symbolic memory addresses and Keystone assembler [69] for generating the ARM Cortex-M patch instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe [70] for the MSP430 patch instructions. Finally, SABRE uses components from the Python library ELFtools [71] to generate the final patched executable. SABRE proof of concept prototype is publicly available at [16].
Table II: Size of SABRE modules
SABRE Static Path Exploit Patch Patch Utilities Total
Module Analysis Verifier Locator Generator Validator
Lines of Code 790 221 1672 1073 941 523 9122
Size (KB) 34.9 8.55 69.5 42.2 39.7 19.2 215.1
6 Evaluation
6.1 Code Size and TCB
Tab. II presents the size of SABRE modules in lines of code and total file size for each SABRE module. Exploit Locator is the largest as it implements the backward traversal and symbolic data flow analysis, the most complex processes of SABRE. Since building a patched CFG is a part of Patch Validator, this support is included in its calculated size, making it the second largest.
6.2 Run-times of SABRE Sub-Modules
@@
Next, we evaluate the time to run each module in SABRE after it has determined that received CFA evidence contains a malicious path. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) [72], we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities (e.g., following a similar methodology to recent work [12]). We then insert them into the example benchmark applications [67] to closely mimic existing MCU-based CVEs [48, 47, 73, 40, 39, 38, 74] to the best of our ability. First, the main function is inspected to determine all function calls. Then, we insert the vulnerability into the last function called by the main. This vulnerability can be used to overwrite $addr_{target}$ to point back to the main function, creating an infinite loop in the program. We then execute the vulnerable programs on two unmodified CFA architectures: a TEE-based approach built atop ARM Cortex-M33 (TRACES [10]) and a custom hardware-based approach built atop TI MSP430 (ACFA [6]). We then collect $\mathcal{C}F_{Log}$ from this execution and pass it to SABRE along with the binary of the modified application (App.elf). Upon receiving $\mathcal{C}F_{Log}$, SABRE executes its sub-modules for further analysis. Fig. 10 shows the run-time for each sub-module.
6.2.1 Static Analysis
@@
The time required to build the CFG via static analysis depends on the total number of branch instructions. Building the CFG for the MSP430 binaries requires more execution time, albeit smaller executables, as shown in Fig. 10(a). This is due to jump destinations in disassembled MSP430 binaries being decoded with offsets, not full addresses [75]. Therefore, MSP430 binaries require additional processing steps to determine jump destinations from the offsets, unlike ARM binaries, which disassemble into full addresses [76]. Generating the CFG from MSP430 binaries required $\approx$66.27 to 70.21ms, whereas ARM binaries required $\approx$9.46 to 17.23ms, depending on the particular application.
6.2.2 Path Validator
@@
The time to verify $\mathcal{C}F_{Log}$ is presented in Fig. 10(b). The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES [10]) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA [6]). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
6.2.3 Exploit Locator
@@
Fig. 10(c) and Fig. 10(d) show the run-time for the Exploit Locator main submodules: the Backwards Traversal and Symbolic Data Flow (DF) Analysis. The Symbolic DF Analysis over $\mathcal{C}F_{Slice}$ incurs the highest run-time from 0.868 to 19.5s. For the OVF example in ARM, $\mathcal{C}F_{Slice}$ of the jfdctint program covers the highest number of instructions (7,750 in total), and thus, it takes the longest time to execute. In MSP430, the aha-compress program (abbreviated compress) has the same characteristic.
@@
Since emulation of memory and registers is not required for Backwards Traversal, its run-time is more efficient in comparison (ranging from $\approx$0.026 to 5.31ms). For a similar reason as described above, Backwards Traversal of UAF in jfdctint requires the longest run-time. The run-time of Backwards Traversal for cover also stands out. The cover application is designed to test the performance while traversing every path in multiple large switch statements. Thus, the run-time slowdown arises due to tracking definitions across the memory stores/loads of each switch case. This effect is especially apparent in ARM binaries, which implement more memory operations for switch statements than MSP430. Overall, the run-time of the Exploit Locator is mainly dominated by the Symbolic DF Analysis.
6.2.4 Patch Generator
Patches are generated by first constructing a patched version of the instructions pertaining to the vulnerability. Then, the ELF executable is updated by adding a new dedicated section with the generated patch code. Times for these steps for MSP430 and ARM Cortex-M33 binaries are shown in Fig. 10(e) and Fig. 10(f), respectively.
@@ 6.2.4 Patch Generator
Overall, the UAF patch is produced faster than the OVF patch due to its simplicity. Generating the patches for ARM Cortex-M33 requires more runtime due to its larger instruction set, differences in the compilation process (i.e., functions in ARM Cortex-M tend to include more loads/stores), and differences in the tooling used to generate the binary for the patch.
6.2.5 Patch Validator
Run-times of each step to verify MSP430 and ARM Cortex-M33 patches are shown in Fig. 10(g), Fig. 10(h), and Fig. 10(i).
@@ 6.2.5 Patch Validator
The time to generate the translated $\mathcal{C}F_{Log}$ depends on the number of new branch instructions introduced as a part of the patch. Consequently, translating $\mathcal{C}F_{Log}$ for the UAF test cases requires negligible runtime since they introduce no new branch instructions (e.g., $<$1.0ms to perform this check, as shown in Fig. 10(g)).
@@
The time to generate the patched CFG depends on the architecture, as described in Sec. 6.2.1. For the UAF case, the runtime is approximately the same since the patch effectively removed one CFG node. The OVF case requires more time to rewrite additional nodes. This difference is more pronounced for MSP430 due to the differences in decoding of jump instructions, as described in Sec. 6.2.1.
@@
Similar to the Exploit Locator, the execution time of the Path Verifier is dominated by the Symbolic DF Analysis (ranging from 0.82s to 14.3s, as shown in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.3). Future work could improve the efficiency of modeling memory accesses in SABRE.
6.3 Patch Sizes
Tab. III shows the increase in binary size (for MSP430 and ARM Cortex-M33) before and after the OVF patch for each application. The baseline (i.e., pre-patched) binary size is larger in ARM Cortex-M33 than in MSP430 due to several reasons. First, compared to the 16-bit MSP430 ISA, ARM Cortex-M is 32-bit and thus requires a larger instruction size. Second, the ARM Cortex-M33 binaries require additional instrumentation by the underlying CFA architecture TRACES [10] for secure generation of run-time evidence. On the other hand, the MSP430 applications are executed alongside a hardware-based CFA architecture (ACFA) that requires no instrumentation and thus incurs no size overhead. Since the baseline size of MSP430 binaries is smaller, their relative post-patched increase is larger (4.58-32.7%). In contrast, the ARM Cortex-M33 patch incurs a smaller relative increase (1.87-3.68%). Since the UAF patch simply replaces a call instruction in place with an equivalent number of nop-s, it incurs no increase in the binary size. However, it does incur an additional data memory overhead at runtime due to the missing free. The data memory overhead depends on the size of the object that is no longer freed.
Table III: Binary size comparison (bytes) for OVF patches
Buffer Overflow
MCU Platform MSP430 ARM Cortex-M33
Application Original Patched Increase (%) Original Patched Increase (%)
libbs 808 1072 32.7% 5212 5404 3.68%
lcdnum 872 1136 30.3% 5692 5880 3.3%
jfdctint 5764 6028 4.58% 7292 7480 2.58%
fibcall 792 1056 33.3% 5308 5496 3.54%
crc32 1236 1500 21.4% 5228 5416 3.6%
cover 3516 3780 7.51% 10268 10460 1.87%
compress 2940 3208 9.12% 6796 6988 2.83%
7 Related Work
7.1 Control Flow Attestation and Auditing
C-FLAT [4] proposed the concept of CFA by leveraging ARM TrustZone to detect and log control flow transfers. In C-FLAT, the binary is instrumented at branching instructions to trap execution in a TrustZone-protected “Secure World”. Once in the Secure World, the branching information is added to a hash chain. At the end of execution, C-FLAT’s hash chain produces a unique value representing the attested program’s control flow path. Inspired by C-FLAT, many follow-on techniques combine instrumentation and TEE support to improve CFA [53, 54, 50, 63].
Several subsequent hardware-based CFA approaches were proposed to remove the requirement of binary instrumentation by presenting custom hardware extensions for detecting control flow transfers [62, 51, 52]. In hardware-based CFA, the hardware extension includes dedicated modules for identifying branch instructions as they execute. The branch information is hashed on-the-fly by extending the MCU with a hardware-based hash engine. Although these techniques remove the requirement for instrumentation, the hardware cost of a fully hardware-based solution is too costly for MCUs.
Tiny-CFA [49] reduces the hardware cost significantly by executing an instrumented binary atop APEX [60], a low hardware-cost PoX architecture. ACFA [6] presents a hybrid (hardware/software co-design) that does not use hash engines, reducing the hardware cost for creating the evidence while still removing the requirement for instrumentation. ACFA’s custom hardware actively interrupts execution to transmit fixed-sized evidence slices to $\sf{\mathcal{V}rf}$. It is the first technique to enable run-time auditing by guaranteeing the eventual delivery of evidence.

### v2

@@ 4.3. Patch Generator
Patching Use-After-Free.
@@
Patching Buffer Overflows.
@@
    Replace the original call to the function from T3 with a trampoline to the patched version.
T1: Estimating Buffer Bounds. Buffer overflows are eliminated via bound checks before memory accesses according to the buffer size. From the previous phase, the Exploit Locator identified the offending instruction to be located at $addr_{Acc}$ and that it causes an overflow onto $base$. This suggests that a buffer exists next to $base$ that must have been overwritten by a loop before the overflow occurs. It also implies that the offending instruction overflowing to $base$ is the same as that used to write data to the buffer.
@@
As depicted in Fig. 8, the buffer’s definition is influenced by -4(r4) at 0xe27e, which is in turn initialized by r11 at 0xe206. Tracking of definitions continues until a known pointer value is reached, as described in Sec. 4.2. In this example, definitions are tracked until 0xe0e4 since the stack pointer (a known pointer value) has been reached. Thus, Exploit Locator saves 0xe0e4 as $addr_{lower}$ (the address defining the lower bound of the exploited buffer) and saves the value of $reg_{Acc}$ (sp in this example) as the lower bound of the overflown buffer at this instruction.
@@
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects and the buffer of interest), narrowing the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by restricting overflows within the buffer’s own stack frame.
T2: Recording the Bounds. After identifying $addr_{lower}$ and $addr_{upper}$, SABRE completes T1 and has estimated the bounds of the buffer that overflows into $addr_{target}$. To complete T2, the values of $reg_{Acc}$ must be saved at both locations. This is completed by placing a trampoline both at $addr_{lower}$ and $addr_{upper}$ to preserve their values in a dedicated register. Fig. 9 depicts the entire patch for buffer overflow vulnerabilities and shows how these instructions are added for recording the bounds.
Figure 9. Example patch for buffer overflow vulnerabilities.
@@
T3: Inserting the Bounds Check Then, Patch Generator proceeds to patch the offending memory instruction by prepending it with instructions that check if the address being accessed is outside the valid boundary. If it is, the offending instruction is skipped. For the patch depicted in Fig. 9, the prepended instructions (highlighted in red) correspond to:
@@
T4: Calling the Safe Function. To realize T4, Patch Generator rewrites the original call instruction that leads to the function containing the buffer overflow. It changes this call instruction to instead call the patched version. This is depicted in Fig. 9 by the instructions highlighted in green.
@@ 4.4. Patch Validator
Before deploying the patched software on $\sf{\mathcal{P}rv}$, it must be validated to ensure its effectiveness against the previously identified attack. To achieve this, this module first builds an updated CFG of the patched binary. Then, it remaps the destination addresses in $\mathcal{C}F_{Slice}$ to their respective addresses in the updated CFG. It also inserts new entries into $\mathcal{C}F_{Slice}$ as necessary to account for the patching behaviors (e.g., a trampoline to/from the patch). Patch Validator then performs symbolic DF analysis over the translated $\mathcal{C}F_{Slice}$, similar to how it is done in Sec. 4.2. When the analysis reaches the end of $\mathcal{C}F_{Slice}$ without $addr_{target}$ being corrupted, Patch Validator considers this patch effective. When the corruption still occurs, the attack was likely caused by another vulnerability. In this case, SABRE generates a report indicating to $\sf{\mathcal{V}rf}$ that further manual analysis is required.
libbslcdnumjfdctintfibcallcrc_32covercompress0020202020404040406060606080808080
(a) Build CFG
libbslcdnumjfdctintfibcallcrc_32covercompress0010101010202020203030303040404040
(b) Verify $\mathcal{C}F_{Log}$
libbslcdnumjfdctintfibcallcrc_32covercompress00222244446666
(c) Backwards Traversal
libbslcdnumjfdctintfibcallcrc_32covercompress05K10K15K20K
(d) Symb. DF Analysis
libbslcdnumjfdctintfibcallcrc_32covercompress00222244446666888810101010
(e) Patch Generation
libbslcdnumjfdctintfibcallcrc_32covercompress00202020204040404060606060
(f) Update ELF
libbslcdnumjfdctintfibcallcrc_32covercompress00111122223333
(g) Translate $\mathcal{C}F_{Log}$
libbslcdnumjfdctintfibcallcrc_32covercompress0020202020404040406060606080808080100100100100
(h) Build Patched CFG
libbslcdnumjfdctintfibcallcrc_32covercompress04K8K12K16K
(i) Symb. Patch Validation
Legend: ARM-UAF: ARM-OVF: MSP-UAF: MSP-OVF:
Figure 10. Run-time of SABRE submodules (ms) to analyze evidence of buffer overflow (OVF) and use-after-free (UAF) vulnerabilities on two MCUs and CFA architectures: hardware-based CFA atop MSP430 (Caulfield et al., 2023) and TEE-based CFA atop ARM Cortex-M33 (Caulfield et al., 2024a).
5. Implementation & Evaluation
@@
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. BEEBs benchmark applications for embedded platforms (Pallister et al., 2013) were used for the evaluation. They were compiled for MSP430 and ARM Cortex-M33 and executed atop unmodified CFA architectures, ACFA (Caulfield et al., 2023) and TRACES (Caulfield et al., 2024a), to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. The MSP430 prototype was configured with 8 KB of application program memory, and the ARM Cortex-M33 prototype was configured with 256 KB of application program memory. We believe that SABRE concepts generalize to several CFA architectures and use ACFA and TRACES as case studies due to their open-source availability.
@@
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy (Team, 2024) for evaluating symbolic expressions and Keystone assembler (Keystone, 2024) for generating ARM Cortex-M instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe (Swiftloke, 2024) for MSP430 instructions. Finally, SABRE uses the Python library ELFtools (Foundation, 2024c) to generate the final patched executable. SABRE proof of concept prototype is publicly available at (Caulfield et al., 2024b).
5.1. Run-times of SABRE Sub-Modules
@@
We evaluate the run-time of each SABRE submodule after a malicious path was detected. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) (Tan et al., 2024), we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities, following a similar methodology to recent related work (Chilese et al., 2024). We then insert them into the example benchmark applications (Pallister et al., 2013) to closely mimic existing CVEs related to control flow hijacking and ROP in MCUs (National Vulnerability Database, 2021a, 2017, 2020a, 2020b, 2020c, b, 2022) to the best of our ability. For consistency across applications, we insert the vulnerability into the last function called by the main. The vulnerability is used to overwrite $addr_{target}$ to return/jump to the main function, causing an infinite loop. We execute the vulnerable programs on two unmodified CFA architectures – a TEE-based approach built atop ARM Cortex-M33 (TRACES (Caulfield et al., 2024a)) and a custom hardware-based approach built atop TI MSP430 (ACFA (Caulfield et al., 2023)) – to collect $\mathcal{C}F_{Log}$-s of the attack, which are used by SABRE with the modified application binary (App.elf) for further analysis. Fig. 10 shows the run-time of each SABRE submodule during the analysis.
@@
Static Analysis. The time required to build the CFG depends on the number of branch instructions. Building the CFG for MSP430 binaries requires more run-time, albeit smaller executables, as shown in Fig. 10(a). This is due to jump destinations in disassembled MSP430 binaries being decoded with offsets, not full addresses (Foundation, 2024b). Therefore, MSP430 binaries require additional processing steps to determine addresses from the offsets, unlike ARM binaries, which disassemble into full addresses (Foundation, 2024a). Generating the CFG from MSP430 binaries required $\approx$66.27 to 70.21ms, whereas ARM binaries required $\approx$9.46 to 17.23ms, depending on the application.
@@
Path Verifier. Fig. 10(b) presents the time to verify $\mathcal{C}F_{Log}$. The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES (Caulfield et al., 2024a)) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA (Caulfield et al., 2023)). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
@@
Exploit Locator. Run-time for Backwards Traversal and Symbolic DF Analysis and are shown in Fig. 10(c) and Fig. 10(d), respectively. The Symbolic DF Analysis incurs the highest run-time from 0.868s to 19.5s. For the OVF example in ARM, $\mathcal{C}F_{Slice}$ of the jfdctint program covers the highest number of instructions (7750 in total), thus having the longest run-time. In MSP430, the aha-compress program (abbreviated compress) has the same characteristic.
@@
Since tracking of memory and registers is not required for Backwards Traversal, its run-time is more efficient in comparison ($\approx$0.026 to 5.31ms). For a similar reason as described above, Backwards Traversal of UAF in jfdctint requires the longest run-time. The cover application is designed to test the performance while traversing every path in multiple large switch statements. Thus, the run-time slowdown arises due to tracking definitions across the memory stores/loads of each switch case. This effect is especially apparent in ARM binaries, which implement more memory operations for switch statements than MSP430. Overall, the run-time of the Exploit Locator is mainly dominated by the Symbolic DF Analysis.
@@
Patch Generator. Overall, the UAF patch is produced faster than the OVF patch due to its simplicity, as shown in Fig. 10(e) and Fig. 10(f). Generating ARM Cortex-M33 patches requires more run-time due to its larger instruction set, differences in the compilation process (i.e., functions in ARM Cortex-M tend to include more loads/stores), and differences in the compiler tools.
@@
Patch Validator. The time to translate the $\mathcal{C}F_{Log}$, shown in Fig. 10(g), depends on the number of new branch instructions introduced as a part of the patch. Therefore, translating $\mathcal{C}F_{Log}$ for the UAF test cases requires negligible run-time (e.g., $<$1.0ms) since it introduces no new branch instructions, whereas the OVF test case incurs time due to new branches introduces with patch.
@@
The time to generate the patched CFG depends on the architecture. For the UAF case, the run-time is approximately the same since the patch effectively removed one CFG node. The OVF case requires more time to rewrite additional nodes. This difference is more pronounced for MSP430 due to decoding of jump instructions, as previously described for Static Analysis.
@@
Similar to the Exploit Locator, the run-time of the Patch Validator is dominated by the Symbolic DF Analysis (from 0.82s to 14.3s in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.2). Future work could improve SABRE’s modeling of memory accesses.
5.2. Patch Sizes
We compare program memory size before and after the patch is installed. As a baseline, the MSP430 binaries consumed from 792 to 5764 bytes (9.90% to 72.1% of the program memory) and the ARM Cortex-M33 binaries consumed from 5.3 KB to 10.3 KB (0.31% to 2.25% of the program memory). The difference in baseline is due to the MSP430 prototype being much more resource-constrained than the ARM Cortex-M33 prototype (see Sec. 5).
After the OVF patch, the MSP430 binaries consumed 1056 to 6028 bytes (13.2% to 75.4% of the program memory) and ARM Cortex-M33 binaries consumed 5.4 KB to 10.4 KB (from 0.41% to 2.35% of the program memory). This represents an average increase of $\approx$264.5 bytes ($\approx$3.31%) and $\approx$189.7 bytes ($\approx$0.10%), respectively.
Since the UAF patch simply replaces a call instruction in place with an equivalent number of nop-s, it incurs no increase in the binary size. However, it incurs additional data memory overhead at run-time due to the missing free. The data memory overhead depends on the size of the object that is no longer freed.
6. Related Work
Control Flow Attestation. C-FLAT (Abera et al., 2016) proposed CFA by instrumenting an application binary at branch instructions with calls that trap execution in an ARM TrustZone-protected “Secure World”, where the branch destination is added to a hash chain representing the current path in a unique digest. Many techniques follow C-FLAT by combining instrumentation and TEE support to improve CFA (Armanuzzaman et al., 2025; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2024a; Yadav and Ganapathy, 2023) or by employing custom hardware extensions to detect/log control flow transfers without requiring instrumentation/TEE support (Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Caulfield et al., 2023).